40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 335     Entry time: Fri Feb 22 14:45:06 2008
Author: steve 
Type: Update 
Category: MOPA 
Subject: laser power levels 

At the beginning of this 1000 days plot shows the laser that was running at 22C head temp
and it was send to LLO

The laser from LHO PA#102 with NPRO#206 were installed at Nov. 29, 2005 @ 49,943 hrs
Now,almost 20,000 hrs later we have 50% less PSL-126MOPA_AMPMON power
Attachment 1: lpower1000d.jpg  98 kB  | Hide | Hide all
lpower1000d.jpg
ELOG V3.1.3-